OpenStudy (anonymous):

convert polar equation to rectangular r=6

OpenStudy (therealmeeeee):

Welcome to OpenStudy!

OpenStudy (therealmeeeee):

By squaring both sides: r^2 = 4^2 ==> r^2 = 16. Then, since r^2 = x^2 + y^2, we see that this equation in rectangular form is: x^2 + y^2 = 16. https://answers.yahoo.com/question/index?qid=20110323100247AA8W9Lm

OpenStudy (therealmeeeee):

If I helped you can close this question